OpenStudy (anonymous):

how do you simplify a rational exponent problem like this? x3/2 y7/3 z9/6

OpenStudy (amistre64):

rational exponents are radicals: x^3/2 = sqrt(x^3) = x sqrt(x)

OpenStudy (anonymous):

im so confused as to how that is typed...

OpenStudy (amistre64):

a thousand monkeys at a thousand typewriters ;)

OpenStudy (amistre64):

the bootom of the fraction part becomes the root of your radical.

OpenStudy (amistre64):

or bottom, depending on the vernacular

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (amistre64):

the answer would be.... take the rational exponents and convert them to radicals :)